# EvaLoop: Assessing LLM Robustness in Programming from a Self-consistency Perspective

### v2

@@ Abstract
Assessing the programming capabilities of Large Language Models (LLMs) is crucial for their effective use in software engineering. Current evaluations, however, predominantly measure the accuracy of generated code on static benchmarks, neglecting the critical aspect of model robustness during programming tasks. While adversarial attacks offer insights on model robustness, their effectiveness is limited and evaluation could be constrained. Current adversarial attack methods for robustness evaluation yield inconsistent results, struggling to provide a unified evaluation across different LLMs. We introduce EvaLoop  [^1] The code is being organized, we will make it public as soon as possible., a novel assessment framework that evaluate the robustness from a self-consistency perspective, i.e., leveraging the natural duality inherent in popular software engineering tasks, e.g., code generation and code summarization. EvaLoop initiates a self-contained feedback loop: an LLM generates output (e.g., code) from an input (e.g., natural language specification), and then use the generated output as the input to produce a new output (e.g., summarizes that code into a new specification). EvaLoop repeats the process to assess the effectiveness of EvaLoop in each loop. This cyclical strategy intrinsically evaluates robustness without rely on any external attack setups, providing a unified metric to evaluate LLMs’ robustness in programming. We evaluate 16 prominent LLMs (e.g., GPT-4.1, O4-mini) on EvaLoop and found that EvaLoop typically induces a 5.01%-19.31% absolute drop in pass@1 performance within ten loops. Intriguingly, robustness does not always align with initial performance (i.e., one-time query); for instance, GPT-3.5-Turbo, despite superior initial code generation compared to DeepSeek-V2, demonstrated lower robustness over repeated evaluation loop. EvaLoop reveals distinct degradation patterns across different models, offering developers a simple-yet-effective assessment of how LLMs maintain functional coherence through multiple transformation loops. This is a critical factor for real-world software development, where requirements and implementations continuously evolve.
@@ 1 Introduction
Figure 1: The generated solutions for MBPP+ No. 472 task by o3-mini. See full results in Appendix D.
@@
  • Empirical Validation and Insights: Through empirical evaluation on 16 popular LLMs, we demonstrate that EvaLoop provides a stable and effective metric for quantifying programming robustness—an LLM’s ability to maintain functionally correct code through iterative modifications. Our findings indicate that applying EvaLoop typically leads to a 5.01%-19.31% absolute decrease in pass@1 performance within ten loops. Intriguingly, we observed that robustness does not necessarily correlate with initial code generation quality; for instance, GPT-3.5-Turbo chatgpt , despite superior initial performance compared to DeepSeek-Coder-V2-Lite-Instruct zhu2024deepseek , exhibited lower robustness (6.38 vs. 6.36) and final performance after evaluation by EvaLoop. Consequently, EvaLoop reveals distinct degradation patterns across LLMs, offering developers a preliminary assessment of how LLMs maintain functional coherence in different evaluation loops. This insight is practical for real-world software development, where requirements and implementations continuously evolve. Furthermore, EvaLoop also offers a novel perspective on the memorization versus comprehension debate in LLMs: the capacity to generate correct solutions over more loops may indicate a deeper comprehension to the prompt, rather than mere recall of training data.
@@ 3 Evaluation
$\text{ASL}=\frac{\sum_{i=1}^{M}i\times n_{i}}{N},$ (1)
where $i$ means the loop index [^2] In code translation, per programming language translation is a loop, $n_{i}$ represents the number of tasks that passed only at loop $i$, and $N$ represents the total number of tasks. To ensure generalizability and representativeness, we conducted an extensive evaluation across 16 state-of-the-art LLMs, encompassing both proprietary closed-source models and open-source alternatives. See details in Appendix A.
@@
Figure 3: Distribution of average sustainable cycles per task in MBPP Plus. The data integrates results from greedy decoding and temperature sampling across 16 popular LLMs (sorted from smallest to largest according to the results of greedy decoding).
@@
Greedy Decoding vs. Temperature. Figure 3 delineates the performance distribution, measured in ASL per task on the MBPP Plus benchmark, across 16 leading LLMs employing either greedy decoding or temperature-controlled stochastic sampling. The analysis reveals a counter-intuitive finding: greedy decoding achieves a marginally superior average performance (6.64 loops) over temperature sampling (6.48 loops). This outcome is particularly noteworthy given that temperature sampling is often the pre-selected decoding mechanism in numerous LLMs (e.g., ChatGPT employs a default temperature of 0.7), primarily valued for its capacity to introduce variability and emulate human-like expression. A plausible explanation for this finding is that the advantages of temperature sampling, such as improved "naturalness" in textual output, are less critical in the domain of code generation. Code inherently demands strict adherence to precise semantic and syntactic rules, a context in which the deterministic, highest-probability path selection of greedy decoding may offer a more consistent advantage over the stochastic nature of temperature sampling. This finding motivates the need for further research into optimal decoding strategy selection specifically tailored for the domain of code generation. Based on the results, we only focus on the results of greedy encoding in the rest of the experiments.
@@
Examining ASL metric reveals that o3-mini achieves the best performance not only on the initial accuracy (80.95% at pass@1) but also consistent robustness (ASL with 7.58) throughout each evaluation loop. While Qwen2.5-Coder-32B exhibits superior performance in the last three evaluation loops, its comparatively modest initial accuracy results in a slightly lower ASL value with 7.42. This underscores how our proposed ASL metric effectively balances performance across all evaluation stages, providing a more comprehensive assessment than single-point measurements. Particularly noteworthy is LLaMA-3.1-8B, which consistently underperforms across all metrics at all evaluation points—especially in ASL, where it scores (below 5) substantially lower than other LLMs. This pronounced disparity highlights the significant challenges some models face in maintaining robustness throughout iterative evaluation loops, reinforcing the value of our framework in identifying models with sustained performance integrity versus those that degrade substantially under realistic usage conditions.
@@
Examining ASL metric reveals another insight into LLMs’ robustness in programming. GPT-4.1 leads with an ASL of 3.42, closely followed by GPT-4.1-mini (3.40) and o3-mini (3.40), signifying not only strong initial accuracy but also remarkable performance consistency across sequential translation tasks, underscoring their robust nature and the superior. In stark contrast, open-source LLMs generally represent a relatively low ASL value, and LLaMA-3.1-8B records the lowest ASL with 2.11, whose trajectory shows a low baseline followed by continuous, pronounced degradation, highlighting difficulties in maintaining functional integrity. ASL is invaluable for a holistic assessment, moving beyond single-point evaluations to capture sustained operational efficacy and distinguish models with enduring performance integrity from those that falter under the cumulative stress of multi-step, cross-lingual code translations.
@@ 5 Conclusion
In this paper, we addressed the critical gap in current evaluations of LLMs’ programming capabilities, which often overlook robustness in favor of initial accuracy. To this, we introduced EvaLoop, a novel framework that assesses LLMs’ robustness in programming from a self-consistency perspective, i.e., leveraging the duality of software engineering tasks, such as code generation and summarization. EvaLoop employs a self-contained feedback loop, iteratively transforming code or its natural language specification until functional degradation occurs, thereby achieving unified evaluation to LLMs’ robustness without relying on external adversarial configurations. Our empirical evaluation across 16 prominent LLMs demonstrated that EvaLoop with framework 1 effectively reveals distinct degradation patterns, typically inducing a 5.01%-19.31% absolute drop in pass@1 performance within ten loops, and importantly, showed that initial programming proficiency does not always correlate with sustained robustness. Hence, EvaLoop provides a more authentic and nuanced analysis of how LLMs maintain functional coherence under iterative stress, offering unique insights for their practical deployment in dynamic software development environments and contributing to the broader discussion on model understanding versus memorization.
@@ Appendix B Limitation
EvaLoop provides significant insights into LLM programming robustness, yet several limitations warrant discussion. Primarily, our assessment relies exclusively on MBPP-derived benchmarks (MBPP Plus and MXEVAL). Despite their enhancements—rigorous quality filtering and expanded test cases—these benchmarks may not fully mirror the diversity of real-world programming tasks. We plan to extend EvaLoop to other benchmark families like HumanEval, APPS, and domain-specific datasets to validate the broader applicability of our findings. Furthermore, computational limitations constrained our evaluation to 16 LLMs, preventing a comprehensive survey of all existing models. Although we carefully selected a mix of leading open and closed-source models based on their MBPP Plus performance and resource accessibility, incorporating additional models, especially novel architectures or specialized coding LLMs, could reveal more nuanced robustness patterns throughout the LLM ecosystem.
@@ Appendix C Experimental Environment
We conducted all experiments on a Ubuntu server with an AMD Ryzen Threadripper PRO 7955WX CPU, 256 GB RAM, and dual NVIDIA A6000 GPUs.
Appendix D Full Loop-Trace Examples for Frameworks 1 & 2
@@
Appendix E Extra Result Exhibition for Framework 2

### v3

@@ Abstract
Assessing the programming capabilities of Large Language Models (LLMs) is crucial for their effective use in software engineering. Current evaluations, however, predominantly measure the accuracy of generated code on static benchmarks, neglecting the critical aspect of model robustness during programming tasks. While adversarial attacks offer insights on model robustness, their effectiveness is limited and evaluation could be constrained. Current adversarial attack methods for robustness evaluation yield inconsistent results, struggling to provide a unified evaluation across different LLMs. We introduce EvaLoop  [^1] The code is being organized, we will make it public as soon as possible., a novel assessment framework that evaluate the robustness from a self-consistency perspective, i.e., leveraging the natural duality inherent in popular software engineering tasks, e.g., code generation and code summarization. EvaLoop initiates a self-contained feedback loop: an LLM generates output (e.g., code) from an input (e.g., natural language specification), and then use the generated output as the input to produce a new output (e.g., summarizes that code into a new specification). EvaLoop repeats the process to assess the effectiveness of EvaLoop in each loop. This cyclical strategy intrinsically evaluates robustness without rely on any external attack setups, providing a unified metric to evaluate LLMs’ robustness in programming. We evaluate 17 prominent LLMs (e.g., GPT-4.1, O4-mini) on EvaLoop and found that EvaLoop typically induces a 5.01%-19.31% absolute drop in pass@1 performance within ten loops. Intriguingly, robustness does not always align with initial performance (i.e., one-time query); for instance, Qwen2.5-Coder-32B-Instruct, despite inferior initial code generation (first five loops) compared to o1, o1-mini, and o4-mini, demonstrated higher robustness over the whole evaluation loop. EvaLoop reveals distinct degradation patterns across different models, offering developers a simple-yet-effective assessment of how LLMs maintain functional coherence through multiple transformation loops. This is a critical factor for real-world software development, where requirements and implementations continuously evolve.
@@ 1 Introduction
Figure 1: The generated solutions for MBPP+ No. 472 task by o3-mini. See full results in Appendix E.
@@
  • Empirical Validation and Insights: Through empirical evaluation on 17 popular LLMs, we demonstrate that EvaLoop provides a stable and effective metric for quantifying programming robustness—an LLM’s ability to maintain functionally correct code through iterative modifications. Our findings indicate that applying EvaLoop typically leads to a 5.01%-19.31% absolute decrease in pass@1 performance within ten loops. Intriguingly, we observed that robustness does not necessarily correlate with initial code generation quality; for instance, Qwen2.5-Coder-32B-Instruct hui2024qwen2 , despite inferior initial performance compared to o1, o1-mini, and o4-mini chatgpt , exhibited higher robustness (7.38 vs. 7.29, 7.26, and 7.32) and final performance when finishing the evaluation. Consequently, EvaLoop reveals distinct degradation patterns across LLMs, offering developers a preliminary assessment of how LLMs maintain functional coherence in different evaluation loops. This insight is practical for real-world software development, where requirements and implementations continuously evolve. Furthermore, EvaLoop also offers a novel perspective on the memorization versus comprehension debate in LLMs: the capacity to generate correct solutions over more loops may indicate a deeper comprehension to the prompt, rather than mere recall of training data.
@@ 3 Evaluation
$\text{ASL}=\frac{\sum_{i=1}^{M}n_{i}\times i^{2}}{MN},$ (1)
where $i$ denotes the loop index [^2] In code translation, per programming language translation is a loop, $M$ represents the total number of evaluation loops, $n_{i}$ represents the number of tasks that passed only at loop $i$, and $N$ represents the total number of tasks. We employ quadratic weighting ($i^{2}$) rather than linear weighting ($i$) to assign progressively higher weights to later evaluation loops, as tasks that maintain functionality through more iterations demonstrate superior model robustness. This quadratic weighting scheme reflects the principle that sustained performance under repeated stress is a stronger indicator of true robustness than early success alone. The normalization by $MN$ scales the results to a range approximately between $\frac{1}{M}$ and $M$, facilitating comparison across different experimental configurations with varying numbers of loops and tasks. To ensure generalizability and representativeness, we conducted an extensive evaluation across 17 state-of-the-art LLMs, encompassing both proprietary closed-source models and open-source alternatives. See details in Appendix A.
@@
Figure 3: Distribution of average sustainable cycles per task in MBPP Plus. The data integrates results from greedy decoding and temperature sampling across 17 popular LLMs (sorted from smallest to largest according to the results of greedy decoding).
@@
Greedy Decoding vs. Temperature. Figure 3 delineates the performance distribution, measured in ASL per task on the MBPP Plus benchmark, across 17 leading LLMs employing either greedy decoding or temperature-controlled stochastic sampling. The analysis reveals a counter-intuitive finding: greedy decoding achieves a marginally superior average performance (6.64 loops) over temperature sampling (6.48 loops). This outcome is particularly noteworthy given that temperature sampling is often the pre-selected decoding mechanism in numerous LLMs (e.g., ChatGPT employs a default temperature of 0.7), primarily valued for its capacity to introduce variability and emulate human-like expression. A plausible explanation for this finding is that the advantages of temperature sampling, such as improved "naturalness" in textual output, are less critical in the domain of code generation. Code inherently demands strict adherence to precise semantic and syntactic rules, a context in which the deterministic, highest-probability path selection of greedy decoding may offer a more consistent advantage over the stochastic nature of temperature sampling. This finding motivates the need for further research into optimal decoding strategy selection specifically tailored for the domain of code generation. Based on the results, we only focus on the results of greedy encoding in the rest of the experiments.
@@
Examining ASL metric reveals that o3-mini achieves the best performance not only on the initial accuracy (80.95% at pass@1) but also consistent robustness (ASL with 7.46) throughout each evaluation loop. While Qwen2.5-Coder-32B exhibits superior performance in the last three evaluation loops, its comparatively low initial accuracy results in a slightly lower ASL value with 7.38 (rank-2). This underscores how our proposed ASL metric effectively balances performance across all evaluation stages, providing a more comprehensive assessment than single-point measurements. Particularly noteworthy is LLaMA-3.1-8B, which consistently underperforms across all metrics at all evaluation points—especially in ASL, where it scores (below 5) substantially lower than other LLMs. This pronounced disparity highlights the significant challenges some models face in maintaining robustness throughout iterative evaluation loops, reinforcing the value of our framework in identifying models with sustained performance integrity versus those that degrade substantially under realistic usage conditions.
@@
Examining ASL metric reveals another insight into LLMs’ robustness in programming. GPT-4.1 and GPT-4.1-mini leads with an ASL of 2.89, closely followed by o3-mini (2.83), signifying not only strong initial accuracy but also remarkable performance consistency across sequential translation tasks, underscoring their robust nature and the superior. In stark contrast, open-source LLMs generally represent a relatively low ASL value, and LLaMA-3.1-8B records the lowest ASL with 1.29, whose trajectory shows a low baseline followed by continuous, pronounced degradation, highlighting difficulties in maintaining functional integrity. ASL is invaluable for a holistic assessment, moving beyond single-point evaluations to capture sustained operational efficacy and distinguish models with enduring performance integrity from those that falter under the cumulative stress of multi-step, cross-lingual code translations.
@@ 5 Conclusion
In this paper, we addressed the critical gap in current evaluations of LLMs’ programming capabilities, which often overlook robustness in favor of initial accuracy. To this, we introduced EvaLoop, a novel framework that assesses LLMs’ robustness in programming from a self-consistency perspective, i.e., leveraging the duality of software engineering tasks, such as code generation and summarization. EvaLoop employs a self-contained feedback loop, iteratively transforming code or its natural language specification until functional degradation occurs, thereby achieving unified evaluation to LLMs’ robustness without relying on external adversarial configurations. Our empirical evaluation across 17 prominent LLMs demonstrated that EvaLoop with framework 1 effectively reveals distinct degradation patterns, typically inducing a 5.01%-19.31% absolute drop in pass@1 performance within ten loops, and importantly, showed that initial programming proficiency does not always correlate with sustained robustness. Hence, EvaLoop provides a more authentic and nuanced analysis of how LLMs maintain functional coherence under iterative stress, offering unique insights for their practical deployment in dynamic software development environments and contributing to the broader discussion on model understanding versus memorization.
@@ Appendix B Limitation
EvaLoop provides significant insights into LLM programming robustness, yet several limitations warrant discussion. Primarily, our assessment relies exclusively on MBPP-derived benchmarks (MBPP Plus and MXEVAL). Despite their enhancements—rigorous quality filtering and expanded test cases—these benchmarks may not fully mirror the diversity of real-world programming tasks. We plan to extend EvaLoop to other benchmark families like HumanEval, APPS, and domain-specific datasets to validate the broader applicability of our findings. Furthermore, computational limitations constrained our evaluation to 17 LLMs, preventing a comprehensive survey of all existing models. Although we carefully selected a mix of leading open and closed-source models based on their MBPP Plus performance and resource accessibility, incorporating additional models, especially novel architectures or specialized coding LLMs, could reveal more nuanced robustness patterns throughout the LLM ecosystem.
@@ Appendix C Experimental Environment
We evaluate most LLMs on an Ubuntu server equipped with an AMD Ryzen Threadripper PRO 7955WX CPU, 256 GB RAM, and dual NVIDIA A6000 GPUs. For larger models exceeding 100 billion parameters, we conduct evaluations on a dedicated deep learning server featuring 8 H200 GPUs and an Intel Xeon Platinum 8470 CPU.
Appendix D Comparison of Linear Weighting and Quadratic Weighting
Table 2: Comparison of ×$\times i$ and ×$\times i^{2}$
ASL (×$\times i$) model name model name ASL (×$\times i^{2}$)
7.585 o3-mini o3-mini 7.457
7.495 o4-mini qw32b 7.385
7.463 o1 gpt-4.1 7.356
7.431 o1-mini o4-mini 7.320
7.418 gpt-4.1 gpt-4.1-mini 7.291
7.418 qw32b o1 7.288
7.349 gpt-4.1-mini dsv2_full 7.267
7.328 dsv2_full o1-mini 7.257
7.119 gpt-4o gpt-4o 7.040
6.913 gpt-4-turbo gpt-4-turbo 6.821
6.471 qw7b qw7b 6.411
6.384 dsv2 dsv2 6.297
6.357 gpt-3.5-turbo gpt-3.5-turbo 6.258
6.013 ds7b ds7b 5.959
5.548 cs22b cs22b 5.485
5.265 oc8b oc8b 5.154
4.053 lm8b lm8b 3.818
The quadratic weighting scheme (×$\times i^{2}$) demonstrates clear advantages over linear weighting (×$\times i$) in capturing true model robustness, as evidenced by the comparison between qw32b and several OpenAI models (gpt-4.1, o1, o1-mini, and o4-mini). As shown in Table 2, under linear weighting, these OpenAI models achieve identical or higher ASL scores than qw32b, suggesting equivalent or superior robustness levels. However, examination of their performance curves (i.e., Figure 4) reveals fundamentally different stability patterns: while the OpenAI models exhibit higher performance in the initial few loops, they demonstrate notable performance decline in subsequent iterations. In contrast, qw32b maintains remarkably consistent performance throughout almost all evaluation loops (after the second loop) with minimal degradation. This discrepancy reveals that linear weighting fails to capture the true robustness differences observed in the performance curves. The quadratic weighting successfully differentiates these behaviors, ranking qw32b higher (7.385) than gpt-4.1 (7.356), o1 (7.288), o1-mini (7.257), and o4-mini (7.320), which accurately reflects qw32b’s superior sustained performance under iterative stress. This demonstrates that the $i^{2}$ weighting effectively amplifies the significance of later-loop performance, providing a more nuanced assessment of model robustness that linear weighting fails to capture when models exhibit similar early performance but divergent long-term stability characteristics.
Appendix E Full Loop-Trace Examples for Frameworks 1 & 2
@@
Appendix F Extra Result Exhibition for Framework 2